A man arranges to pay off a debt of 3600 euros by 40 annual instalments which form an arithmetic series.when 30 of the instalments are paid he dies leaving a third of the debt unpaid;find the value of first instalment.

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Abhishek

a + (a+d) + (a+2d)+......+(a+39d)  = 3600          (1)

(a+30d) +.....+ (a+39d) = 3600/3 = 1200            (2)

use ∑n = n(n+1)/2

solving we get

a=51, d=2

first instalment= 51 euro

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..